# Diffusion Model with Perceptual Loss

### 26 Feb 2024 v4

@@ 2.1 Diffusion Models Generate Bad Samples
Let’s first analyze why diffusion models in their raw form generate bad samples. As illustrated in Figure 0(a), consider a simple distribution where the images always contain a solid circle at arbitrary locations. Given the model only a few training samples (top row), we want the model to generate a new sample from this distribution (bottom left), but the actual generation can be far out of the distribution (bottom right).
From a non-biased standpoint, the model’s generation of two circles is equally plausible given the limited training samples. When observations are sparse in a distribution or a sub-region of a distribution, there is room for ambiguities.
Another factor that can cause the issue is model capacity. Even when more training samples are given, the model must have enough capacity to learn and deduce that the underlying distribution most likely always contains one circle only.
Figure 1: Illustration of the current diffusion model problem. Diffusion models learn a plausible pixel distribution from the training samples but it does not align with the distribution of real images.
As illustrated in Figure 0(b), similar problems exist for actual images. The diffusion model tries to learn a plausible pixel distribution from finite training samples, but it does not always align with the intended distribution of photorealistic images.
Existing works have focused on expanding the dataset [32, 4], increasing the model capacity [24, 4, 14, 22, 29, 27], simplifying the ODE trajectory [18, 17, 13], and improving the sampler [36, 19, 20]. However, almost all works use the mean squared error (MSE) loss in training and no study has explored the use of perceptual loss.
@@ 2.3 Perception in Classifier Guidance
However, we argue that the surprising effectiveness of classifier guidance also partially originates from it being an implicit form of perceptual guidance. Specifically, the classifier network, i.e. CLIP [26], is a trained semantic neural network that can provide perceptual guidance. The guidance gradients are high on more perceptually important areas, such as the shape and contour, and low on the imperceivable areas. The diffusion model is implicitly guided toward what the perceptual classifier considers as more probable, resulting in more perceptually photorealistic samples.
@@ 2.4 Perception in Classifier-Free Guidance
Classifier-free guidance (CFG) [11] can also be viewed from the lens of perceptual guidance. The diffusion model itself is a neural network that learns high-level semantic features, and classifier-free guidance exploits the perceptual knowledge from the model itself.
Classifier-free guidance is defined as:
$x_{cfg}=x_{\oslash}+w(x_{c}-x_{\oslash})$ (1)
We first query the model unconditionally $x_{\oslash}$, then query the model conditionally $x_{c}$. We subtract them to get the “difference”, then accentuate this difference by a large factor $w$.
This “difference” is in fact the high-level semantic perception that is the most essential about the target prompt. The difference is large in more perceptually important areas and small in unimportant areas. In this implicit way, classifier-free guidance is a form of perceptual guidance as well.

### 06 Mar 2024 v5

@@ 2.1 Diffusion Models Generate Bad Samples
Let’s first elucidate why diffusion models in their raw form generate bad samples. As shown in Figure 1, given finite training samples, the underlying data distribution is ambiguous. The maximum likelihood estimation (MLE) is a distribution that assigns even probability only to the observed samples and zero everywhere else. An ideal score model will learn this ground-truth probability flow to always produce observed samples and generate no new data. In practice, neural networks do not overfit to this flow exactly, allowing generalization and thus the generation of new data.
Figure 1: The underlying data distribution is ambiguous given finite training data.
Figure 2: Diffusion models learn a plausible pixel distribution from the training samples but it does not align with the distribution of real images.
However, the learned distributions may not match the real underlying distribution. As illustrated in Figure 1(a), consider a simple distribution where the images always contain a solid circle at arbitrary locations. Given limited training samples (top row), we want to generate a new sample from this distribution (bottom left), but the actual generation can be far out of the distribution (bottom right). This problems exist for actual images as illustrated in Figure 1(b). The diffusion model tries to learn a plausible pixel distribution from finite training samples, but it does not align with the intended distribution of photorealistic images.
The ground-truth probability flow is determined by the dataset and forward diffusion function alone, but the learned distribution is determined by the model capacity and loss function. Existing works have focused on expanding the dataset [32, 4], improving the model architecture [24, 4, 14, 22, 29, 27], simplifying the ODE trajectory [18, 17, 13], and improving the sampler [36, 19, 20]. However, almost all works use the mean squared error (MSE) loss in training and no study has explored the use of perceptual loss.
@@ 2.3 Perception in Classifier Guidance
However, we argue that the surprising effectiveness of classifier guidance also partially originates from it being a form of perceptual guidance. Specifically, the classifier network, i.e. CLIP [26], is a trained perceptual neural network that can provide perceptual guidance. The classifier gradients guide toward perceptually more probable images as a precondition to better text alignment. This results in the generation of more photorealistic images.
@@ 2.4 Perception in Classifier-Free Guidance
Classifier-free guidance (CFG) [11] finds that an implicit classifier can be derived from using Bayes’ rule: $p(c|x_{t})\propto p(x_{t}|c)/p(x_{t})$ and the score model itself can be used to provide guidance $\nabla\log p(c|x_{t})\propto s(x_{t}|c)-s(x_{t})$. Therefore, classifier-free guidance can also be viewed from the lens of perceptual guidance.